To estimate the mean score μ of those who took the Medical College Admission Test on your campus, you will obtain the scores of an SRS of students. From published information you know that the scores are approximately Normal with standard deviation about 6.2. You want your sample mean x to estimate ,1 with an error of no more than 0.8 point in either direction. (a) What standard deviation must x have so that 99.7% of all samples give an x within 0.8 point of μ? Use the 68-95-99.7 rule. (Enter your answer rounded to four decimal places.) standard deviation of - (b) How large an SRS do you need in order to reduce the standard deviation of rounded to the nearest whole number) to the value you found? (Enter your answer SRS size-
Answer #1

Ans:

a)99.7% of the data falls within 3 standard deviations of mean,so

Margin of error=0.8

3*=0.8

=0.8/3=0.2667

b)

=6.2/sqrt(n)=0.2667

Sample size required,n=(6.2/0.2667)^2=541